Thank you all for helping, here is my question.
Explain how to complet the square for a binomial. Use x^2 + 6x to illustrate your explaination.

Answer by ankor@dixie-net.com(12701)

:
Use x^2 + 6x to illustrate your explanation.
:
Write it:
x^2 + 6x + ___ = 0
:
Choose a value that will complete the square, divide the coefficient of x by 2 and square it.
In this problem: 6/2 = 3, squared = 9, so we add 9 to both sides of the equation
x^2 + 6x + 9 = + 9
:
(x + 3)^2 = 9
:
x + 3 = +/-Sqrt(9)
:
x + 3 = +/-3
:
x = -3 + 3
x = 0
and
x = -3 -3
x = -6